### v3

# Oxygen Vacancies in Niobium Pentoxide as a Source of Two-Level System Losses in Superconducting Niobium

Abstract
We identify a major source of quantum decoherence in three-dimensional superconducting radio-frequency (SRF) resonators and two-dimensional transmon qubits composed of oxidized niobium: oxygen vacancies in the niobium pentoxide which drive two-level system (TLS) losses. By probing the effect of sequential in situ vacuum baking treatments on the RF performance of bulk Nb SRF resonators and on the oxide structure of a representative Nb sample using time-of-flight secondary ion mass spectrometry (ToF-SIMS), we find a non-monotonic evolution of cavity quality factor $Q_{0}$ which correlates with the interplay of Nb2O5 vacancy generation and oxide thickness reduction. We localize this effect to the oxide itself and present the insignificant role of diffused interstitial oxygen in the underlying Nb by regrowing a new oxide via wet oxidation which reveals a mitigation of aggravated TLS losses. We hypothesize that such vacancies in the pentoxide serve as magnetic impurities and are a source of TLS-driven RF loss.
preprint: APS/123-QED
I Introduction
The realization of a computationally useful superconducting quantum computer requires the maximization of quantum coherence lifetimes to unlock quantum memory, high-fidelity gates, and better entanglement among qubits [1]. In superconducting systems, quantum coherence is limited by dissipative decay channels present in both the linear resonator and the transmon qubit that contains the nonlinear Josephson junction necessary to create uniquely addressable energy levels. Current decoherence candidates include quasi-particles [2, 3], radiation-induced losses [4, 5], and dielectric losses such as two-level systems (TLS) [6], to name a few. Dielectric loss arises from each material and interface used in the construction of superconducting qubits and dictates the total decay rate of the system according to
$\Gamma_{1}=\frac{1}{T_{1}}=\frac{\omega}{Q_{0}}=\omega\sum_{i}\frac{F_{i}}{Q_{i}}+\Gamma_{0},$ (1)
where $T_{1}$ is the photon lifetime, $\omega$ is angular frequency, $Q_{0}$ is the quality factor, $F_{i}$ and $Q_{i}$ are the filling and quality factors of the $i^{th}$ loss channel, and $\Gamma_{0}$ captures losses driven by non-dielectric mechanisms [7, 8]. The quality factor is inversely related to the dielectric loss tangent $Q_{i}=1/\text{tan}\delta_{i}$. As a result, decay rate minimization requires the mitigation of every loss channel down to tolerable levels. While straightforward for a single interface, this task is difficult in complex multi-layer systems like those used in modern qubit fabrication designs. Many such designs utilize thin films of aluminum and niobium atop dielectric or semiconducting substrates, leading to multiple potentially deleterious materials and interfaces.
Recent work on three-dimensional (3-D) niobium superconducting radio-frequency (SRF) resonators in the low GHz frequency range has highlighted the amorphous native niobium oxide as a dramatic source of TLS loss, implying an inherent limitation in qubits fabricated with oxidized niobium [9, 10]. By removing the native niobium oxide via in situ vacuum baking at 340-450C for several hours, a ten-fold improvement in quality factor $Q_{0}$ was observed, demonstrating intra-resonator photon lifetimes of $T_{1}\approx$ 2 s, a 200-fold increase over the previous state-of-the-art [11]. By fitting temperature dependent $Q_{0}$ data of bulk niobium resonators with the standard TLS model [12, 13], Romanenko et al. found the dielectric loss tangent of the native niobium oxide at 5 GHz to be $\delta_{0}$ = 0.08. This value is in agreement with values obtained from niobium-on-silicon lumped element resonators post sequential oxidation [14] and niobium-on-silicon coplanar-waveguides post sequential etching [15]. This consistency between dramatically different architectures demonstrates that observations of the Nb-air interface made in simplified, single interface systems such as 3-D Nb SRF resonators are transferable to more complex systems such as those used in transmon qubits, allowing for a one-to-one correspondence between thin film and bulk material measurements.
While there exists an empirical mitigation strategy to minimize their deleterious effect in superconducting quantum systems, the precise origin and nature of the TLS losses within the niobium oxide is not yet fully understood. A complete picture of the underlying mechanisms would aid in the further development of mitigation strategies that may instead be stable to air. There exist a few possible origins for these TLS losses. One possibility stems from the conducting or magnetic nature of the NbOx compounds [10]. Another lies in the theory that interstitial oxygen in the niobium lattice could provide tunneling sites for hydrogen [16]. A third possibility involves oxygen vacancies within the amorphous pentoxide acting as magnetic impurities [17, 18] and driving impedance, as supported by X-ray absorption spectroscopy studies [19] and density functional theory calculations [19, 20]. Here, we present results that support the role of the latter in driving TLS losses in oxidized niobium.
One avenue for identifying the origins of loss in the amorphous native Nb oxide is by studying its evolution with in situ baking, allowing for the investigation of the role of various oxides on $Q_{0}$. There exists extensive literature on this topic due to its relevance in high $Q_{0}$ and high electric field applications of Nb SRF resonators in particle accelerators. It is well known that native niobium oxides exist in a graded structure composed of Nb2O5 (insulator [21]), NbO2 (semiconductor [22]), NbO (metallic [23]), and the metal substrate [21, 24, 25]. In situ vacuum baking dissolves the dominant pentoxide and grows the suboxides, reducing according to Nb2O5 $\rightarrow$ NbO2 $\rightarrow$ NbO [24, 26, 27, 28]. Ma et al. showed that in situ vacuum baking at 150C for 40 hours reduced the pentoxide to NbO2; continued baking at 280C for 8 hours showed further reduction to metallic NbxO ($\sim$2) [26]. Further studies by Ma et al. revealed the reduction of the pentoxide to NbO2 and NbO post vacuum baking at 250C for several hours [28]. These results are corroborated by the synchrotron measurements of Delheusy et al. [27] and ARXPS measurements by Dacca et al. [23]. Moreover, at the expense of dissolving the oxide, in situ vacuum baking drives inward oxygen diffusion from the oxide over the oxide/metal interface rather than into the vacuum [26, 29, 30, 27, 31]. Subsequent reoxidation studies via exposure to ambient air shows that a partially dissolved oxide grows at the oxide/metal interface and is morphologically equivalent to the native grown oxide without prior heat treatment [28].
In this paper, by investigating the role of vacuum baking on niobium SRF resonators which contain a $\sim$5 nm full wet-grown native amorphous oxide, we find that vacancies in the niobium pentoxide host TLS which limit the performance of 3-D Nb resonators and 2-D transmon qubits which utilize oxidized Nb. Modest in situ vacuum baking treatments (150C-200C) for durations as short as 25 minutes aggravate TLS-like losses that are eventually suppressed with increasing bake time. We recreate the in situ vacuum baking treatments using time-of-flight secondary ion mass spectrometry (ToF-SIMS) on a Nb SRF resonator cutout and confirm that the oxide reduces considerably, suggesting an increase in the number of oxygen vacancies in the Nb2O5 layer. Continued in situ baking shows that the total number of oxygen vacancies in the pentoxide decreases as the layer is gradually reduced into suboxides. As such, the non-monotonic evolution of $Q_{0}$ with successive baking is due to the interplay of oxygen vacancy generation in the Nb2O5 layer, which host TLS, and subsequent pentoxide dissolution. Finally, we consider one potential model in which these oxygen vacancies drive TLS, namely, that they host magnetic impurities.
II Experimental Method
For our studies, we used three 1.3 GHz TESLA-shaped [32] niobium single-cell SRF resonators that have undergone a bulk 120 $\mu$m removal from the inner RF surface via electropolishing, followed by 800C degassing, and an additional 40 $\mu$m removal from the inner surface via electropolishing [33]. The resonators were then exposed to air and high pressure rinsed with ultra-pure water, forming a $\sim$5 nm amorphous native oxide. Afterwards, the resonators were assembled for testing and evacuated to a vacuum level of approximately 1E-5 Torr. Typically, resonators tested at this point are called electropolished (EP) resonators. The resonators then underwent sequential testing and treatment to various low temperature vacuum bakes [34] for increasing durations while actively pumping to maintain the vacuum level within the resonators. Such treatments have been shown to dissolve the native niobium oxide and introduce oxygen interstitial into the underlying niobium [26, 29, 30, 27, 35, 31].
One resonator, TE1AES019, was first subjected to an in situ low temperature vacuum bake at 90C for 384 hours followed by sequential in situ baking steps at 200C for increasing durations. For the final treatment, the resonator was baked at 340C for 5 hours to ensure complete dissolution of the niobium pentoxide [10]. The resonator was RF tested after each step and maintained vacuum throughout the entire course of the study.
The second resonator, TE1RI003, was used to isolate losses to the dynamics of the oxide itself. After its baseline test post electropolishing, the cavity underwent in situ low temperature baking at 200C for 1 hour followed by further testing. The cavity was then vented to clean room air and left to oxidize for a full week. This drove the re-formation of the native niobium oxide but this time with a larger concentration of subsurface O interstitial present compared to the baseline measurement post EP. Once complete, TE1RI003 was once again re-evacuated and tested.
The third resonator, TE1AES021, was used to probe baking treatments that more closely resembled those typically utilized in transmon qubit fabrication and to further identify the source of additional RF loss. It underwent sequential in situ vacuum baking treatments at 150C with resonator RF measurements performed after each step. After the RF test post baking at 150C for a net duration of 15 hours and 25 minutes, resonator TE1AES021 was subjected to an ex situ vacuum bake at 200C for 19.5 hours. This involved first venting the resonator to atmosphere followed by vacuum baking. Afterwards, the resonator was once again high pressure rinsed, reforming the amorphous $\sim$5 nm native oxide, again increasing the concentration of subsurface interstitial O compared to the baseline EP treatment. For the final treatment, the resonator underwent an HF acid rinse, which dissolved the native oxide, and upon subsequent high pressure rinsing, reformed a new native oxide at the expense of consuming approximately 2 nm of niobium [36].
[…]
III Results
III.1 Resonator RF measurements
The upper panel in Fig. 1 presents data acquired on resonator TE1AES019 after an initial 90$\times$ 384 hour in situ vacuum bake followed by sequential in situ vacuum bakes at various temperatures and durations. Shown also for reference is data from Romanenko and Schuster taken on an EP resonator which contains a full 5 nm native wet-grown oxide and serves as our baseline test [9]. For all tests, we observe the expected saturation of the quality factor at electric fields $<$ 0.01 MV/m (i.e. photon numbers $<$ 2E18) that is characteristic of two-level systems [9, 10, 6]. We find that the 90$\times$ 384 hours $+$ 200$\times$ 1 hour in situ vacuum baking treatment produces a resonator with a low field quality factor of 9E9, roughly three times lower than the baseline test. An additional 200$\times$ 5 hour in situ vacuum bake produces identical results at fields below 0.1 MV/m. After a second round of 200$\times$ 5 hour vacuum bake, a slight improvement in $Q_{0}$ up to 1.2E10 occurs. For the final treatment, we subjected the resonator to a 340$\times$ 5 hour vacuum baking treatment which is known to fully dissolve the native Nb2O5 layer [10, 40]. The resulting measurements display a $Q_{0}$ at low fields of greater than 7E10, the highest achieved in this study, and in line with previous work [10, 40]. Overall, these tests show a non-monotonic dependence of resonator $Q_{0}$ with bake duration.
Figure 1: Measurements of quality factor vs electric field at the center of the cavity at temperatures of 1.4 K-1.6 K for the 1.3 GHz 3-D Nb resonators a) TE1AES019, b) TE1RI003, and c) TE1AES021 after sequential baking and chemical treatments. Upper axes plot the converted intra-resonator photon number. For reference, the gray left pointed triangles in a) and c) depict data obtained on an EP resonator which contained a full 5 nm native wet grown oxide from Romanenko and Schuster [9]. Solid black lines show fits with Eq. 2.
Fig. 1b) instead plots the results of cavity TE1RI003. The baseline test post EP yields an expected low field $Q_{0}$ of 2.5E10. Post 200$\times$ 1 hour in situ baking, we observe a reduction of this value by a factor of approximately three, in agreement with that observed in Fig. 1a). Subsequent re-exposure to clean room air for one week and further testing showed a mitigation of aggravated TLS losses and a return to EP baseline performance. These results highlight the role of oxide dynamics in driving cavity performance.
The lower panel in Fig. 1 depicts data taken on resonator TE1AES021 post various baking and chemical treatments. Surprisingly, we find that a modest in situ vacuum bake at 150C for 25 minutes reduces the resonator $Q_{0}$ by a factor of two when compared to the baseline EP test from Romanenko and Schuster. Subjecting the resonator to an additional round of in situ vacuum baking at 150C for 15 hours produces a low field $Q_{0}$ of 9E9, identical to the results obtained on TE1AES019 and TE1RI003 post 200$\times$ 1 hour in situ baking. The results post ex situ bake at 200C for 19.5 hours, for which we re-iterate received subsequent high pressure rinsing with ultra-pure water, yield a $Q_{0}$ that is similar in value to the baseline EP at fields $<$ 4E-3 MV/m. Subsequent HF acid rinsing and high pressure water rinsing reveals a $Q_{0}$ vs $E_{acc}$ curve that is nearly identical to the baseline EP data.
We fit the data in Fig. 1 with the standard TLS model with a quasi-particle term built in. The electric field dependence of a niobium SRF resonator at low fields may be modelled as a single interface system (metal/air) containing TLS according to
$\frac{1}{Q_{0}}=\frac{F\delta_{0}}{\big{(}1+\big{(}\frac{E_{acc}}{E_{c}}\big{)}^{2}\big{)}^{\beta}}+\frac{1}{Q_{qp}}$ (2)
where $\beta$ is a fitting parameter, $E_{c}$ is the characteristic saturation field, $\delta_{0}$ is the dielectric loss, and $Q_{qp}$ is the quasi-particle contribution to RF loss [13, 9]. Fig. 2 plots the product of the filling factor and dielectric loss tangent for each curve presented in Fig. 1. We choose to present this convolved product as it is agnostic to the precise values of the filling factor and loss tangent and gives a qualitative measure of the TLS contribution to RF loss. Resonator TE1AES019 shown in the upper panel of Fig. 2 displays a non-monotonic dependence of the TLS-driven losses with in situ vacuum baking, achieving as low as 3.63E-12 after the final round of treatment, consistent with previous results [10]. TE1RI003 presented in the middle panel of Fig. 2 further suggests that in situ vacuum baking increases TLS-driven RF losses while reoxidation in dry clean room air for one week returns them to baseline EP values. The results of TE1AES021 in the lower panel further support this observation and suggest that a fully oxidized surface, whether dry or wet grown, contains less two-level system loss than a partially dissolved oxide.
Figure 2: Fitted $F\delta_{0}$ values and fit error of a) TE1AES019, b) TE1RI003, and c) TE1AES021 using the data presented in Fig. 1. The 90$\times$ 384 hours in situ vacuum bake that precedes the sequential in situ vacuum baking treatments for resonator TE1AES019 has been omitted from the horizontal axis for brevity. Most error bars are smaller than the symbol size.
We mention that the characteristic saturation field extracted from Fig. 1 remains roughly constant, as evidenced by the increase in $Q_{0}$ at 0.1-0.2 MV/m for all investigated cavities and temperatures. This supports the theory that a single loss mechanism remains dominant throughout these studies: two-level systems located within the oxide.
III.2 ToF-SIMS studies
As the amorphous niobium oxide is known to host TLS, it is likely that its dynamics dictate the performance evolution observed in Figs. 1 and 2. To investigate this possibility, we recreated the in situ baking treatments presented in Fig. 1a) on an electropolished niobium SRF resonator cutout in the preparation chamber associated with a ToF-SIMS setup. This allowed us to perform treatments and measure changes in the surface chemistry without re-exposing to ambient conditions. We used a liquid bismuth ion Bi+ beam to perform secondary ion mass measurements while sputtering with a 500 eV cesium ion Cs+ gun to obtain depth profiles of various species. The sample was sequentially vacuum baked in steps of 200C followed by cool down to and measurement at room temperature in situ. The final bake was performed at 340C for 5 hours. Three depth profiles were obtained at different spots on the sample after each step of treatment to ensure sample uniformity. All measurements were performed in a vacuum better than 1E-10 Torr. Fig. 3 presents the resulting averaged Nb2O5-, O-/Nb-, and H-/Nb- signals as a function of depth into the samples.
Figure 3: Averaged ToF-SIMS depth profiles of the Nb2O5-, O-/Nb-, and H-/Nb- signals acquired on an as-received electropolished resonator cutout subjected to sequential rounds of in situ vacuum baking. The Nb2O5- signals are scaled relative to one another such that the values at 10 nm are identical. Nb2O5 thickness $t$ is defined as the depth at which counts drop below 10, as denoted by the dashed line. Durations presented in the legend represent integrated bake times.
The as-received electropolished resonator cutout displays a fully intact 5 nm thick native pentoxide. After the first round of in situ baking at 200C for 1 hour, we find that this Nb2O5 layer is reduced slightly, extending to a thickness of 4 nm. Continued baking shows further reduction in pentoxide thickness and is eliminated by the 340$\times$ 5 hour bake. These results confirm the gradual dissolution of the native Nb2O5 layer with sequential in situ vacuum baking at 200C and the subsequent inward diffusion of oxygen over the oxide/metal interface [26, 29, 30, 27, 31, 35]. Coincident with the dissolution of the oxide is the decrease of the H- signal within it. We note, however, that the H- signal in the niobium metal below the oxide remains unchanged, suggesting that the concentration of hydrogen in the niobium metal remains constant throughout this study.
As discussed above, oxygen vacancies in the niobium pentoxide may serve as a potential source of TLS-driven losses in niobium-based resonators; to investigate if such vacancies correlate with the evolution of our cavity performance, we extract out the number of missing oxygen atoms within the pentoxide layer for each treatment presented in Fig. 3. This is done by first scaling the total oxygen counts by the pentoxide thickness to obtain the expected oxygen counts per unit thickness $\sigma$ for the as-received, EP baseline sample. Second, for each annealing step, we calculate the expected oxygen counts by scaling $\sigma$ with the measured pentoxide thickness $t$. We then calculate the measured oxygen counts $o$ from the O-/Nb- signal in Fig. 3 by integrating from the surface to depth $t$. Lastly, we take the difference between the expected and actual measured oxygen counts such that we have $\sigma t-o$. The result yields the missing oxygen counts in the pentoxide which is proportional to the total oxygen vacancies in the sample. We note that this analysis is not applicable to the results post 340$\times$ 5 hour in situ baking as the pentoxide is eliminated, making it difficult to quantify the number of missing counts in the absent layer. The results presented in Fig. 4 show a non-monotonic evolution with subsequent baking treatments. Such behavior could be explained by considering an interplay of oxygen vacancy generation and overall oxide thickness reduction. At first, baking produces many vacancies while the overall thickness decreases only slightly. However, as baking continues, the oxide thickness decreases dramatically, and the total number of vacancies within the oxide begins to decrease.
Figure 4: Averaged missing oxygen counts in the niobium oxide obtained using data presented in Fig. 3. These values are proportional to the number of oxygen vacancies in the oxide; see text for details. Error bars depict standard deviation.
IV Discussion
The non-monotonic increase and subsequent decrease in TLS-driven RF losses combined with the non-monotonic evolution of the number of missing oxygen counts suggests the role of oxygen vacancies in the pentoxide as a source of TLS. Indeed, previous tunneling and XPS measurements on cavity-grade niobium baked at various temperatures and durations show that, when compared to an unbaked niobium sample, the 250$\times$ 2 hr bake not only causes Nb2O5 to reduce but also yields an increased level of subgap states which is likely driven by magnetic moments formed due to the presence of oxygen vacancies within the oxide [18]. Moreover, other work has shown that these magnetic moments increase with oxygen vacancy concentration [17]. As such, in situ vacuum baking may promote the proliferation of oxygen vacancies that drive TLS-like losses in 3-D niobium resonators. The aggravation and subsequent mitigation of TLS-driven losses post in situ vacuum baking is likely due to the interplay of oxygen vacancy generation in the Nb2O5 layer and overall pentoxide dissolution.
This hypothesis is further corroborated by the fact that treatments which re-expose the resonator to air and reform the native oxide show nearly identical TLS-driven losses when compared to the baseline EP test. Moreover, despite the fact that TE1RI003 post re-oxidation and TE1AES021 post the additional HF acid rinse contained a larger concentration of interstitial oxygen in the underlying niobium material, both gave nearly identical performance to that of the baseline EP values in Figs. 1 and 2. This localizes the TLS losses to the oxide and asserts the non-role of interstitial oxygen in niobium at these fields and temperatures. This further supports the conclusions made in [9].
We find that hydrogen is likely not responsible for the variations in performance that we observe here. From Fig. 3, we find that the hydrogen concentration at the surface decreases monotonically. Moreover, the hydrogen in the niobium metal does not change appreciably with the baking. Neither of these experimental observations can be used to explain the non-monotonic evolution of the RF performance.
[…]
Figure 5: Surface resistance introduced by TLS for resonator TE1AES019 post 90$\times$ 384 hour + 200$\times$ 1 hour shown in the upper panel in Fig. 1.
While the present work focuses on the effect of vacuum baking on 3-D niobium SRF resonator performance in the quantum regime, the observations made here are applicable to transmon qubits where niobium oxides are present. As many qubit fabrication processes perform vacuum baking treatments after the deposition and oxidation of niobium thin films, there exists the potential of generating a magnetic pentoxide which may further limit transmon qubit quantum coherence times.
V Conclusion
In summary, we studied the effect of low temperature vacuum baking on the quality factor of three 1.3 GHz 3-D niobium SRF resonators and found evidence which suggests that oxygen vacancies in the niobium pentoxide are a potential source of TLS losses that limit performance in both Nb resonators and transmon qubits which utilize oxidized Nb. In situ vacuum baking produces an aggravation and subsequent mitigation of these losses due to the interplay of oxygen vacancy generation and oxide dissolution.
VI Acknowledgments
This material is based upon work supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Superconducting Quantum Materials and Systems Center (SQMS) under contract number DE-AC02-07CH11359.